# SPRec: Leveraging Self-Play to Debias Preference Alignment for Large Language Model-based Recommendations

### v1

@@ 3.3. Evaluating Biases and Unfairness through Distribution Alignment Perspective
When aligning user preferences, LLMs may inadvertently learn biased or unfair outcomes. To assess the bias and fairness issues in LRSs, a mainstream perspective is to formulate the problems as a mismatch distribution problem (Dai et al., 2024). Specifically, let $R$ denote the ground-truth user preference (e.g., an item list), following the distribution $P(R)$, and let $\hat{R}$ represent the model-predicted preferences, drawn from $P(\hat{R})$. Bias or unfairness is then quantified by the mismatch between these two distributions:
(5) $P(R)\neq P(\hat{R}).$
@@
To operationalize this framework, we adopt the method outlined by Jiang et al. (2024), where $P(R)$, the target user history preference distribution, is approximated using the category distribution derived from offline training data. Furthermore, we also employ their proposed metrics, MGU@$K$ and DGU@$K$, to systematically measure the degree of mismatch in our experiments.
4. Method
In this section, we first point out the severe popularity bias due to the nature of DPO, then we analyze how we can address this by utilizing the self-play philosophy. Then we detail the proposed SPRec architecture.
4.1. Problem of DPO: Amplify Popular Bias
@@
As we mentioned DPO is effective and efficient in aligning LLMs with the user preferences, it has a problem of suppressing the negative samples too much, which will result in severe popularity bias in the recommendation task. In recommendation tasks, the input and positive samples $(x,y_{w})$ are derived from logged interactions in offline data, while negative samples $y_{l}$ are drawn from non-interacted items. Given an input $x$, the probability $p_{\mathcal{D}}(y|x)$ represents the popularity of item $y$ in the dataset, and $q_{\mathcal{D}}(y|x)$ denotes the probability of the same item $y$ being selected as a negative sample. Using these definitions, the DPO loss in Eq. (4) can be rewritten as:
@@
In most recommendation setting, the probability of an item $y$ being a negative sample satisfies a uniform distribution (Bai et al., 2024; Chen et al., 2024b), i.e., $q_{\mathcal{D}}(y|x)=\mathcal{U}=\frac{1}{|\mathcal{I}|}$, where $\mathcal{I}$ is the all item candidates. Furthermore, in most preference alignment scenarios using the DPO-based method, the parameter $\beta$ is constrained to the range $10<\beta<$ 1. Under these conditions, the DPO loss naturally encourages the policy to prioritize popular items with higher probabilities $p_{\mathcal{D}}(y|x)$, thereby introducing a notable popularity bias.
In the extreme case where $\beta\to 0$, the optimal policy under the DPO loss degenerates to recommending only the most popular items with the highest probabilities $p_{\mathcal{D}}(y|x)$, effectively excluding less popular but potentially valuable items from consideration.
Remark: This result is a byproduct of DPO’s loss function. Unlike the forward KL-divergence $\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$ used in the SFT loss in Eq. (3.1), DPO optimizes the reverse KL-divergence $\mathbb{D}_{KL}(\pi_{\theta}(y|x),\pi_{\mathrm{ref}}(y|x))$. Forward KL-divergence is known for its mass-covering property, which encourages learning an average behavior and is less sensitive to subtle differences in the preference distribution (as illustrated in Fig. 2(a)) (Sun and van der Schaar, 2024). In contrast, the reverse KL-divergence used in DPO promotes mode-seeking behavior (Wang et al., 2023a; Omura et al., 2024), guiding the model to focus on the “peaks” of the distribution (Fig. 2(b)).
@@ 4.2. Solution: Suppress Biases through Self-Play
Since the DPO loss inherently causes the policy $\pi_{\theta}$ to learn sharp ”peaks”, leading to bias, an intuitive solution is to directly suppress these learned peaks. To address this, we utilize a self-play framework, dubbed SPDPO, which iteratively alternates between policy learning and bias suppression. Specifically, in the $(t+1)$-th iteration, negative samples are drawn from the model’s predictive distribution $\pi_{\theta_{t}}(\cdot|x)$ at iteration $t$, resulting in the following learning paradigm:
@@ 4.3. Architecture of SPRec
The iteration will repeat $T$ times in an epoch. The self-play philosophy can utilized in any LLM tuning-based LRSs. Next, we will explain the necessity of the SFT step, which is not considered in original self-play fine-tuning work (Chen et al., 2024a).
@@
In fact, combining SFT and DPO has been shown to be an effective practice in recent research and open-sourced LLM models. For example, each iteration of post-training for Llama 3 includes an SFT stage followed by a DPO stage (Dubey et al., 2024). Similarly, Pang et al. (2024) demonstrate the effectiveness of an iterative preference optimization algorithm useing a modified DPO loss with an additional negative log-likelihood (NLL) term, which mirrors the SFT loss defined in Eq. (1).
Table 1. Statistics of Datasets.
Datasets #Items #Interactions #Sequences
MovieLens 10,682 10,000,054 9,301,274
Goodreads 4,058 160,398 6,031
CDs_and_Vinyl 13,078 185,855 21,347
Steam 32,094 178,961 29,876
@@ 5. Experiments
Table 2. Overall performance comparison of SPRec (green) with SFT-based (brown) and DPO-based (blue) methods. Best results are in bold, sub-optimal results are underlined. Metrics with $\uparrow$ indicate higher is better, while $\downarrow$ indicates lower is better.
Dataset Model DivRatio@1$\uparrow$ DivRatio@5$\uparrow$ ORRatio@1$\downarrow$ ORRatio@5$\downarrow$ MGU@1$\downarrow$ MGU@5$\downarrow$ HR@5$\uparrow$ NDCG@5$\uparrow$
SASRec 0.0133 0.0057 0.6588 0.4778 0.0463 0.0439 0.0240 0.0155
@@
RW 0.2821 0.2646 0.2648 0.0971 0.0432 0.0305 0.0112 0.0103
$\text{D}^{3}$ 0.3135 0.2968 0.2672 0.0748 0.0320 0.0252 0.0103 0.0088
@@
[…]
$\text{D}^{3}$ 0.1580 0.1414 0.4560 0.1896 0.0511 0.0320 0.0730 0.0718
[…]
SDPO 0.2140 0.1978 0.3270 0.0756 0.0448 0.0320 0.0810 0.0804
RosePO 0.2060 0.1868 0.3670 0.1080 0.0562 0.0349 0.0760 0.0750
[…]
5.1. Experimental Setup
5.1.1. Datasets
We conducted extensive experiments on four real-world datasets: MovieLens [^1] https://grouplens.org/datasets/movielens/ (Harper and Konstan, 2015), Steam [^2] https://cseweb.ucsd.edu/j̃mcauley/datasets.html#amazon_reviews, Goodreads [^3] https://mengtingwan.github.io/data/goodreads, and the CDs and Vinyl category of the Amazon Review Dataset [^4] http://jmcauley.ucsd.edu/data/amazon/index_2014.html. Following the data processing approach in (Chen et al., 2024b; Bao et al., 2024), interaction sequences with fewer than 10 entries were excluded. The datasets were then split chronologically into training, validation, and test sets in an 8:1:1 ratio, ensuring mutual exclusivity and preventing data leakage. The statistics of the datasets are summarized in Table 1.
Table 3. Ablation Experiment Results. SPRec-RN represents the training process with negative samples randomly sampled, while SPRec-noDPO and SPRec-noSFT denote SPRec without DPO and SFT, respectively.
Dataset Model DivRatio@1$\uparrow$ DivRatio@5$\uparrow$ ORRatio@1$\downarrow$ ORRatio@5$\downarrow$ MGU@1$\downarrow$ MGU@5$\downarrow$ HR@5$\uparrow$ NDCG@5$\uparrow$
SPRec-noSFT 0.3020 0.2486 0.0837 0.0224 0.0198 0.0188 0.0184 0.0149
[…]
SPRec-noDPO 0.2220 0.2006 0.3910 0.1884 0.0620 0.0294 0.0790 0.0776
[…]
5.1.2. Evaluation Setting
To leverage the strengths of LLMs in generative recommendation tasks, we prompt the LLM to generate a predicted item based on the input history sequence. Then, following the procedures in BIGRec (Bao et al., 2023a), we calculate scores and rankings for the entire item space and ground our predicted item to an exact item in the dataset.
5.1.3. Metrics.
We evaluate the model using four key metrics on a randomly sampled set of 1,000 test cases per iteration. Accuracy is measured using Normalized Discounted Cumulative Gain (NDCG@$k$) and Hit Ratio (HR@$k$) and calculated per result before averaging. Diversity is assessed with DivRatio@$k$, representing the proportion of unique items in the top-$k$ recommendations. Over-recommendation is quantified by ORRatio@$k$, indicating the proportion of the top-$k$ results consisting of the three most frequently recommended items. Fairness is evaluated using MGU@$k$ (Jiang et al., 2024), which captures category-level discrepancies between the model’s top-$k$ recommendations and the user’s historical sequences.
5.1.4. Baseline
For traditional recommendation models, we select SASRec (Kang and McAuley, 2018), a widely used baseline employing a sequential method with a self-attention mechanism. For LLM-based models, we consider several baselines. (1) For SFT-based methods, BIGRec (Bao et al., 2023a) serves as an instruction-tuning LLM framework for sequential recommendations and forms the foundation for SPRec. Re-weighting (RW) (Jiang et al., 2024) improves fairness in BIGRec by balancing recommendations across categories through dataset-based training weights. Debiasing-Diversifying Decoding ($\textbf{D}^{3}$) (Bao et al., 2024) enhances diversity in BIGRec using a decoding strategy guided by SASRec. (2) For DPO-based models, DMPO (Bai et al., 2024) introduces DPO into LRSs by sampling multiple negative items as rejected responses, while Softmax-DPO (SDPO) (Chen et al., 2024b) follows a similar approach but incorporates a softmax loss over multiple negative samples. Finally, RosePO (Liao et al., 2024a) is a preference optimization framework that combines negative sampling strategies and personalized uncertainty to achieve fairness, unbiasedness, and robustness. The implementation details are listed in Appendix B.
5.2. Overall Performance Comparison (RQ1)
The experimental results are shown in Table 5, leading to the following observations.
5.2.1. Limitations of SFT-based Methods
Fine-tuning LLMs with instruction-based methods results in recommendations heavily favoring popular items, leading to a lack of diversity. For example, in the Goodreads dataset, the DivRatio@1 of BIGRec is only 0.142, meaning the model provides just 14 distinct recommendations per 100 tasks. Similarly, in the Steam dataset, BIGRec’s ORRatio@1 reaches 0.391, with over 39% of recommendations concentrated on the 3 most popular items. These findings highlight that relying solely on SFT introduces severe biases, significantly overexposing certain popular items.
5.2.2. Limitations of DPO-based Methods
For DPO methods using random sampling, such as SDPO and DMPO, while multiple negative samples improve recommendation accuracy, they perform poorly on diversity and fairness metrics. In the Goodreads and MovieLens datasets, SDPO and DMPO exhibit significantly higher ORRatio values than BIGRec, supporting the conclusion in Section 4.1 that DPO amplifies popularity bias. On the CDs and Steam datasets, although ORRatio improves, diversity metrics remain largely unchanged, indicating that the model favors moderately popular items but fails to effectively recommend new ones. In summary, existing DPO-based methods not only fail to address fairness issues in LRS but may exacerbate popularity bias.
5.2.3. Superiority of SPRec
SPRec also surpasses most baseline models on DivRatio and ORRatio across the majority of datasets, showcasing its superior ability to balance recommendation distributions. For fairness, SPRec achieves the best MGU@1 on all datasets except Steam. Moreover, as shown in Fig. 3, SPRec alleviates category-level unfairness on the MovieLens dataset, achieving the best results in 7 out of 8 categories, further underscoring its effectiveness in improving fairness.
@@ 5.3.1. SPRec without SFT
As shown in Table 3, SPRec-noSFT achieves the lowest recommendation accuracy across all datasets except Steam. This indicates that, during the self-play process, the model’s excessive focus on fairness compromises its accuracy. In the MovieLens dataset (Fig. 4), the absence of the SFT stage leads to a steady decline in recommendation accuracy (NDCG@1) throughout training. These findings highlight the critical role of SFT in maintaining SPRec’s recommendation performance.

### v3

@@ 3.3. Evaluating Bias via Distribution Alignment
When aligning user preferences, LLMs may inadvertently learn biased or unfair outcomes. To assess the bias and fairness issues in LRSs, a mainstream perspective is to formulate the problems as a mismatch distribution problem (Dai et al., 2024). Specifically, let $R$ denote the ground-truth user preference (e.g., an item list), following the distribution $P(R)$, and let $\hat{R}$ represent the model-predicted preferences, drawn from $P(\hat{R})$. Bias or unfairness is then quantified by the mismatch between these two distributions: $P(R)\neq P(\hat{R})$.
@@
To apply this framework, we follow Jiang et al. (2024), approximating $P(R)$ using the category distribution from offline training data. We further employ their MGU metric to systematically measure the degree of mismatch in our experiments.
4. Problem of DPO: Amplify Popularity Bias
We present an empirical analysis to demonstrate how DPO exacerbates popularity bias in LRSs, followed by a theoretical examination of the underlying mechanisms driving this phenomenon.
Figure 2. Distribution of cold-start recommendation results. Group 0: least popular, group 4: most popular.
Figure 3. Illustration of SFT, DPO, and SPRec in LLM-based recommendations. (a) SFT generates mass-covering results but retains inherent biases. (b) DPO amplifies these biases by over-representing certain items. (c) SPRec mitigates bias through self-play, leveraging model outputs as negative samples to achieve balanced recommendations.
4.1. Empirical Analysis
To systematically evaluate how DPO amplifies recommendation bias, we design a cold-start recommendation task using two widely-used benchmark datasets: MovieLens and Goodreads [^1] Dataset details are provided in Section 6.1.1. In this task, the LLM generates recommendations without access to user interaction history. We randomly sample 100 items and partition them into five groups based on interaction probabilities, ensuring balanced popularity levels. Positive samples are drawn accordingly, while negative samples for DPO training are randomly selected. The training and validation sets contain 4096 and 512 samples, respectively. After training, we analyze the distribution of recommendations across the five groups.
Fig. 2 shows the proportion of recommendations allocated to each group before and after training, where group 0 represents the least popular items and group 4 the most popular. The results reveal three key insights: (1) SFT introduces a slight bias toward group 4; (2) DPO significantly amplifies this bias, causing recommendations to concentrate almost entirely on the most popular items; and (3) our proposed method, SPRec (introduced in detail later), effectively mitigates the bias amplification induced by DPO.
4.2. Theoretical Analysis
@@
In recommendation tasks, the input and positive samples $(x,y_{w})$ are derived from logged interactions in offline data, while negative samples $y_{l}$ are drawn from non-interacted items. Given an input $x$, the probability $p_{\mathcal{D}}(y|x)$ represents the conditional popularity of item $y$ in the dataset, and $q_{\mathcal{D}}(y|x)$ denotes the probability of the same item $y$ being selected as a negative sample. Using these definitions, the DPO loss in Eq. (3) can be rewritten as:
@@
In most recommendation settings, negative samples are uniformly distributed (Bai et al., 2024; Chen et al., 2024b), i.e., $q_{\mathcal{D}}(y|x)=\mathcal{U}=\frac{1}{|\mathcal{I}|}$, where $\mathcal{I}$ is the set of all candidate items. Additionally, in typical DPO-based preference alignment scenarios, $\beta$ is constrained to $10<\beta<$ 1. Under these conditions, the DPO loss inherently biases the model toward popular items with higher $p_{\mathcal{D}}(y|x)$, exacerbating popularity bias. In the extreme case where $\beta\to 0$, the optimal policy collapses to recommending only the most popular items, effectively disregarding less frequent but potentially valuable recommendations.
Remark: This result is a byproduct of DPO’s loss function. Unlike the forward KL-divergence $\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$ used in the SFT loss in Eq. (3.1), DPO optimizes the reverse KL-divergence $\mathbb{D}_{KL}(\pi_{\theta}(y|x),\pi_{\mathrm{ref}}(y|x))$. Forward KL-divergence is known for its mass-covering property, which encourages learning an average behavior and is less sensitive to subtle differences in the preference distribution (as illustrated in Fig. 3(a)) (Sun and van der Schaar, 2024). In contrast, the reverse KL-divergence used in DPO promotes mode-seeking behavior (Wang et al., 2024a; Omura et al., 2024), guiding the model to focus on the “peaks” of the distribution (Fig. 3(b)).
@@
5. Method
We present how to address the popularity bias in DPO by utilizing the self-play philosophy. Then we detail the proposed SPRec architecture.
@@ 5.1. Solution: Suppress Biases through Self-Play
Since the DPO loss inherently causes the policy $\pi_{\theta}$ to learn sharp “peaks”, leading to bias, an intuitive solution is to directly suppress these learned peaks. To address this, we utilize a self-play framework, dubbed SPDPO, which iteratively alternates between policy learning and bias suppression. Specifically, in the $(t+1)$-th iteration, negative samples are drawn from the model’s predictive distribution $\pi_{\theta_{t}}(\cdot|x)$ at iteration $t$, resulting in the following learning paradigm:
@@ 5.2. Architecture of SPRec
This process repeats for $T$ iterations per epoch. The self-play mechanism is adaptable to any LLM-based recommender system. To ensure comparability with existing DPO-based recommenders (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024), we can extend from a single to multiple negative samples, with results analyzed in experiments.
@@
In fact, combining SFT and DPO has been shown to be an effective practice in recent research and open-sourced LLM models. For example, each iteration of post-training for Llama 3 includes an SFT stage followed by a DPO stage (Dubey et al., 2024). Similarly, Pang et al. (2024) demonstrate the effectiveness of an iterative preference optimization algorithm using a modified DPO loss with an additional negative log-likelihood (NLL) term, which mirrors the SFT loss defined in Eq. (1).
@@ 6. Experiments
Table 1. Overall performance comparison of SPRec (green), SFT-based (brown), and DPO-based (blue) methods. Best results are bold, sub-optimal ones underlined. $\uparrow$ indicates higher is better, while $\downarrow$ indicates lower is better.
@@
RW 0.2821 0.2648 0.0432 0.0112 0.0103
$\text{D}^{3}$ 0.3135 0.2672 0.0320 0.0103 0.0088
[…]
SDPO 0.3218 0.2118 0.0380 0.0110 0.0106
[…]
SASRec 0.0010 1.0000 0.1094 0.0660 0.0379
[…]
$\text{D}^{3}$ 0.1580 0.4560 0.0511 0.0730 0.0718
[…]
SDPO 0.2080 0.3510 0.0475 0.0820 0.0810
RosePO 0.2310 0.3160 0.0499 0.0820 0.0805
[…]
6.1. Experimental Setup
6.1.1. Datasets
We conducted extensive experiments on four real-world datasets: MovieLens [^2] https://grouplens.org/datasets/movielens/, Steam [^3] https://cseweb.ucsd.edu/j̃mcauley/datasets.html#amazon_reviews, Goodreads [^4] https://mengtingwan.github.io/data/goodreads, and the CDs and Vinyl category of the Amazon Review Dataset [^5] http://jmcauley.ucsd.edu/data/amazon/index_2014.html. Additional details about the datasets are provided in the Appendix B. Following the data processing approach in (Chen et al., 2024b; Bao et al., 2024), interaction sequences with fewer than 10 entries were excluded. The datasets were then split chronologically into training, validation, and test sets in an 8:1:1 ratio, ensuring mutual exclusivity and preventing data leakage. To ensure comparability across different LLM-based methods, we further sampled 4,096 interactions from each dataset’s training set as the training samples for all methods, 512 interactions from the validation set, and 1,000 interactions from the test set.
[…]
6.1.2. Evaluation Setting
To leverage the strengths of LLMs in generative recommendation tasks, we prompt the LLM to generate a predicted item based on the input history sequence. Then, following the procedures in BIGRec (Bao et al., 2025), we calculate scores and rankings for the entire item space and ground our predicted item to an exact item in the dataset.
6.1.3. Metrics.
We evaluate the model on 1,000 randomly sampled test cases per iteration using four key metrics. Accuracy is measured by NDCG@5 and HR@5, averaged across results. Diversity is assessed via DivRatio, representing the proportion of unique recommendations. Over-recommendation is quantified by ORRatio, indicating the proportion of results dominated by the three most frequently recommended items. Fairness is evaluated using MGU (Jiang et al., 2024), capturing category-level discrepancies between recommendations and user history.
6.1.4. Baseline
For traditional recommendation models, we select SASRec (Kang and McAuley, 2018), a widely used baseline employing a sequential method with a self-attention mechanism. For LLM-based models, we consider several baselines. (1) For SFT-based methods, BIGRec (Bao et al., 2025) serves as an instruction-tuning LLM framework for sequential recommendations and forms the foundation for SPRec. Re-weighting (RW) (Jiang et al., 2024) improves fairness in BIGRec by balancing recommendations across categories through dataset-based training weights. Debiasing-Diversifying Decoding ($\textbf{D}^{3}$) (Bao et al., 2024) enhances diversity in BIGRec using a decoding strategy guided by SASRec. (2) For DPO-based models, DMPO (Bai et al., 2024) introduces DPO into LRSs by sampling multiple negative items as rejected responses, while Softmax-DPO (SDPO) (Chen et al., 2024b) follows a similar approach but incorporates a softmax loss over multiple negative samples. Finally, RosePO (Liao et al., 2024) is a preference optimization framework that combines negative sampling strategies and personalized uncertainty to achieve fairness, unbiasedness, and robustness. The implementation details are listed in Appendix C.
6.2. Overall Performance Comparison (RQ1)
The experimental results are presented in Table 6, leading to the following observations. The non-LLM baseline, SASRec, performs poorly with the given training size, which is expected as SASRec requires large datasets to achieve effective fitting. In this study, we primarily focus on LLM-based methods, and SASRec’s results are included only for reference and as the assistant model for $\text{D}^{3}$ during the decoding stage.
6.2.1. Limitations of SFT-based Methods
Fine-tuning LLMs with instruction-based methods results in recommendations heavily favoring popular items, leading to a lack of diversity. For example, in the Goodreads dataset, the DivRatio of BIGRec is only 0.142, meaning the model provides just 14 distinct recommendations per 100 tasks. Similarly, in the Steam dataset, BIGRec’s ORRatio reaches 0.391, with over 39% of recommendations concentrated on the 3 most popular items. These findings highlight that relying solely on SFT introduces severe biases, significantly overexposing certain popular items.
6.2.2. Limitations of DPO-based Methods
For DPO methods using random sampling, such as SDPO and DMPO, while multiple negative samples improve recommendation accuracy, they perform poorly on diversity and fairness metrics. On the Goodreads and MovieLens datasets, SDPO and DMPO have minimal impact on DivRatio and ORRatio and may even degrade model performance. On the CDs and Steam datasets, although ORRatio decreases, diversity metrics remain largely unchanged, suggesting that the model favors moderately popular items but fails to effectively recommend new ones. In contrast, RosePO performs well on the CD dataset due to its negative sampling strategy based on semantic information. However, this approach heavily relies on the semantic characteristics of the dataset’s structure, resulting in relatively poor performance on other datasets and limiting its generalizability for debiasing.
In summary, existing DPO-based methods fail to address fairness issues in LRS.
6.2.3. Superiority of SPRec
SPRec also surpasses all baseline models on DivRatio and ORRatio, showcasing its superior ability to balance recommendation distributions. For fairness, SPRec achieved the highest MGU scores on the MovieLens and Goodreads datasets, and the second-highest on the Steam and CD datasets. Moreover, as shown in Fig. 4, SPRec alleviates category-level unfairness on the MovieLens dataset, achieving the best results in 7 out of 8 categories, further underscoring its effectiveness in improving fairness.
Table 2. Ablation results. “with RN” for random negative samples, “w/o” for without specific components.
[…]
w/o SFT 0.3020 0.0837 0.0198 0.0184 0.0149
[…]
with RN 0.2194 0.2224 0.0544 0.0286 0.0230
[…]
with RN 0.1380 0.3380 0.0188 0.0420 0.0310
[…]
w/o SFT 0.3381 0.2373 0.0216 0.0132 0.0126
@@ 6.2.3. Superiority of SPRec
Steam SPRec 0.2930 0.2560 0.0367 0.0910 0.0893
@@ 6.3.1. SPRec without SFT
As shown in Table 2, SPRec w/o SFT achieves the lowest recommendation accuracy across all datasets except Steam. This indicates that, during the self-play process, the model’s excessive focus on fairness compromises its accuracy. In the MovieLens dataset (Fig. 5), the absence of the SFT stage leads to a steady decline in recommendation accuracy (NDCG) throughout training. These findings highlight the critical role of SFT in maintaining SPRec’s recommendation performance.